the perimeter df a standard basketball court is 288ft.find the dimentions of the basketball court if it's length is 44ft.longer than it's width.
Answer by orca(409)

Let x be the width.
Then the length is x + 44.
As the perimeter of a rectangle is p = 2*length + 2*width, we have
2(x+44)+2x = 288
Solve the equation for x, we have
x+44 + x = 144 (divide both sides by 2)
2x + 44 = 144
2x = 100
x = 50
So the width is 50, and the length is x + 44 = 94.